### v1

@@ Abstract
The radio access network (RAN) is a critical component of modern telecom infrastructure, currently undergoing significant transformation towards disaggregated and open architectures. These advancements are pivotal for integrating intelligent, data-driven applications aimed at enhancing network reliability and operational autonomy through the introduction of cognition capabilities, exemplified by the set of enhancements proposed by the emerging Open radio access network (O-RAN) standards. Despite its potential, the nascent nature of O-RAN technology presents challenges, primarily due to the absence of mature operational standards. This complicates the management of data and applications, particularly in integrating with traditional network management and operational support systems. Divergent vendor-specific design approaches further hinder migration and limit solution reusability. Addressing the skills gap in telecom business-oriented engineering is crucial for the effective deployment of O-RAN and the development of robust data-driven applications. To address these challenges, Boldyn Networks, a global Neutral Host provider, has implemented a novel cloud-native data analytics platform. This platform underwent rigorous testing in real-world scenarios of using advanced artificial intelligence (AI) techniques, significantly improving operational efficiency, and enhancing customer experience. Implementation involved adopting development operations (DevOps) practices, leveraging data lakehouse architectures tailored for AI applications, and employing sophisticated data engineering strategies. The platform successfully addresses connectivity challenges inherent in offshore windfarm deployments using long short-term memory (LSTM) Models for anomaly detection of the connectivity, providing detailed insights into its specialized architecture developed for this purpose.
@@ 1 Introduction
To address these escalating demands, innovative enhancements in network design have emerged. A recent progress in this arena is the advent of open radio access network (O-RAN) technology. O-RAN aims to disaggregate the monolithic, single-vendor RAN reducing infrastructure costs and paving the way for network programmability, ultimately leading to autonomous network operations by leveraging native-supported in-network artificial intelligence (AI) techniques, aiming to streamline the complexities of designing, delivering, managing and operating private networks. It establishes a new framework of standards and principles for wireless networking, emphasizing open standards, interfaces, functions, and interoperability to foster greater market competition.
O-RAN’s primary objective is to reduce vendor lock-in, enhance flexibility, and develop network cognitive functions by leveraging data to optimize network performance and improve its resilience, to achieve cost efficiency, particularly in managing diverse heterogeneous networks. Its key advantages lie in software-defined network (SDN) technologies and virtualized network functions (VNF), which not only slash deployment costs but also enable network programmability for autonomous management. This, in turn, reduces operational complexity and optimizes performance [1].
@@
On the other hand, O-RAN is still a relatively new approach. It comes with many realistic challenges in implementation, such as interoperability with legacy network management systems, data integration issues, immaturity in data processing platforms to produce data-driven applications, management, and other technical complexities. In addition to these challenges, there is a shortage of experienced engineers and an increased number of engineering roles with specific skill sets that are necessary to boost this transformation in RAN architecture.
There is a wide range of applications required in the O-RAN system to improve its functionality, such as predictive maintenance and anomaly detection, energy efficiency optimization, automated network configuration and healing, enhanced quality of service (QoS) and traffic management, enhanced user admission control, dynamic RAN slicing [2, 3, 4, 5]. AI and machine learning (ML) approaches are usually considered as the main tools to tackle these challenges [6]. The combination of programmability and AI in O-RAN, from the implementation of xApps and rApps using the interfaces offered by RAN intelligent controllers (RICs), leads to the automation of network management tasks and makes real-time, data-driven decisions [7]. The architectural innovations of O-RAN allow network operators to integrate AI algorithms into the O-RAN network, enabling the use of AI for tasks such as network optimization, troubleshooting, and other complex business problems. By automating these tasks, AI can be used to improve network performance, enhance the customer experience, and reduce costs.
However, developing solutions in O-RAN involves numerous challenges and complexities. These include obtaining a large amount of reliable training data from the network, managing and monitoring AI models for execution and inference, and regulating and implementing update mechanisms for the AI models deployed in the operational support system (OSS) stack.
In theory, the data and model aspects of these challenges can be partially addressed through machine learning operations (MLOps) [8]. However, in the context of O-RAN, the challenges are more complex due to multi-purpose applications and multi-platform issues stemming from the multi-vendor nature of O-RAN. The data sources for AI models are vast and intricate, often originating from multiple systems as shown in Fig. 1. Consequently, relying on a traditional data solution provided by a single vendor is nearly impossible in the multi-vendor environment of the mobile network operator industry.
@@
Figure 1: Multi-vendor ORAN network setup with Vendor-A in light blue, Vendor-B in green, and Vendor-Z in orange. In addition to other systems, equipment, and user devices that are working within the network and produce important data.
This underscores the need for a comprehensive platform and methodology to apply analytics and AI solutions within O-RAN. Such an approach is essential to effectively address these challenges and organize the efforts involved in developing data-driven solutions. These challenges present opportunities for network operators to design and implement a platform capable of tackling these issues whilst being able to proactively test and deploy AI models in O-RAN, to enhance the network’s adaptability and efficiency.
In this paper, we introduce a novel cloud-native, open data-driven platform for O-RAN to address these challenges and streamline the integration of AI applications into the O-RAN management stack. This platform leverages cutting-edge engineering technologies and concepts such as data lakehouse [9], DevOps [10], and MLOps. The design of this platform also considers the growing number of private network deployments, and the increased complexities of these networks and use cases, ensuring future scalability and the capacity to handle the vast amounts of data generated within the network. The contributions of this paper can be summarized as follows:
@@
    To the best of the authors’ knowledge, this is the first holistic cloud-native platform for multi-vendor system integration and services management proposed for O-RAN.
@@
    This cloud-native platform is tightly aligned with the O-RAN architecture for potential AI model implementation and integration.
@@
    This platform fully automates involved infrastructure, setup, data, and AI pipelines using DevOps and GitOps technologies, significantly reducing operational workload and streamlining the development cycle, which enables better collaboration between business owners and developers, leading to more efficient resource utilization.
  4. 4.
    The paper presents an existing near real-time business problem related to connectivity to be resolved on an offshore mobile network that uses O-RAN technology and provides a solution for it using anomaly detection with a long-short term memory (LSTM) model in the proposed data platform. This is the first AI-based solution targeting a use case of an offshore mobile network built using O-RAN technology.
  5. 5.
    This paper provides a new cloud-native open data platform architecture that will support multi-vendor O-RAN designs. It also presents the used method of work in the AI lifecycle and other data-driven applications to tackle the lack of specialized human resources and standardization when deploying AI in O-RAN.
The remaining sections of this paper are constructed as follows: Sec. 2 presents the background of O-RAN and AI-enabled intelligent networks. Sec. 3 elaborates on the proposed cloud-native open data platform. The anomaly detection use case leveraging AI technique is detailed in Sec. 4. Then, in Sec. 5, we discuss the potential problems of the raised platform and the plan for its future development. Lastly, the conclusions of this paper in Sec. 6.
2 Preliminaries
In this section, we present the background information regarding the O-RAN technique and its AI applications.
2.1 O-RAN
The RAN is a critical component of a typical mobile communication network, enabling UE to connect to the core network, which then delivers services to users. The evolution of wireless communication systems from 1G to 5G highlights increasing modularity and virtualization of network functionalities.
Key advancements in RAN architecture include distributed RAN (D-RAN), centralized (or Cloud) RAN (C-RAN), and virtual RAN (vRAN). The distinctions among these architectures are detailed in [11].
@@ 2.2 The motivation of deploying AI in O-RAN
  • Reducing Complexity: O-RAN networks have a more complex, disaggregated architecture compared to traditional RAN, making manual management and optimization more challenging. Also, building traditional applications that process the data will be challenging. AI algorithms can automate and optimize these processes. it also compensates for the shortage of skilled engineers to manage such novel networks [15].
@@
    • Improved Network Performance: AI-based algorithms can be used to optimize network performance by dynamically allocating network resources and adjusting network parameters based on real-time network conditions such as autonomous QoE and QoS resource optimization [4, 15].

### v2

@@ Abstract
The Radio Access Network (RAN) is a critical component of modern telecommunications infrastructure, currently evolving towards disaggregated and open architectures. These advancements are pivotal for integrating intelligent, data-driven applications aimed at enhancing network reliability and operational autonomy through the introduction of cognitive capabilities, as exemplified by the emerging Open Radio Access Network (O-RAN) standards. Despite its potential, the nascent nature of O-RAN technology presents challenges, primarily due to the absence of mature operational standards. This complicates the management of data and intelligent applications, particularly when integrating with traditional network management and operational support systems. Divergent vendor-specific design approaches further hinder migration and limit solution reusability. These challenges are compounded by a skills gap in telecommunications business-oriented engineering, which remains a key barrier to effective O-RAN deployment and intelligent application development. To address these challenges, Boldyn Networks developed a novel cloud-native data analytics platform, specifically designed to support scalable AI integration within O-RAN deployments. This platform underwent rigorous testing in real-world scenarios, and applied advanced Artificial Intelligence (AI) techniques to improve operational efficiency and customer experience. Implementation involved adopting Development Operations (DevOps) practices, leveraging data lakehouse architectures tailored for AI applications, and employing sophisticated data engineering strategies. The platform successfully addresses connectivity challenges inherent in real-world offshore windfarm deployments using Long Short-Term Memory (LSTM) models for anomaly detection in network connectivity. After integrating the LSTM models into the network control, more than 90 percent of connectivity issues were reduced in runtime. This marks a step toward autonomous, self-organising, and self-healing networks.
@@ 1 Introduction
To address these escalating demands, innovative enhancements in network design have emerged. A recent progress in this arena is the advent of open radio access network (O-RAN) technology. O-RAN aims to disaggregate the monolithic, single-vendor RAN reducing infrastructure costs and vendor lock-in, and paving the way for network programmability, ultimately leading to autonomous network operations by leveraging data and native-supported in-network artificial intelligence (AI) techniques, aiming to streamline the complexities of designing, delivering, managing and operating private networks. It establishes a new framework of standards and principles for wireless networking, emphasizing open standards, interfaces, functions, and interoperability to foster greater market competition [1].
A wide range of applications are being investigated in O-RAN due to their critical role in improving system functionality, such as predictive maintenance and anomaly detection, energy efficiency optimization, automated network configuration and healing, improved quality of service (QoS) and traffic management, enhanced user admission control, dynamic RAN slicing [2, 3, 4, 5]. AI and machine learning (ML) approaches are considered the main tools that can address these challenges [6].
@@
However, O-RAN is still a relatively new approach and developing solutions in it comes with many realistic implementation challenges, such as interoperability with legacy network management systems, big data problem, data integration issues, the need for sophisticated pipelines for data acquisition and processing to produce AI- and data-driven applications, as well as management and other technical complexities. In addition to these challenges, there is a shortage of experienced engineers and a growing need for engineering roles with specific skill sets that are essential to support the transformation of the O-RAN architecture.
A clear example of these challenges was observed during the deployment of a live offshore private network supporting windfarm vessels in the North Sea. The network, based on O-RAN principles, operated in a complex environment shaped by changing weather conditions, vessel mobility, and the wide area of coverage in the sea. These factors made connectivity management more difficult and resulted in recurring connectivity issues. Existing tools offered limited visibility and control, making it difficult to identify and address the root causes. Addressing this required more than just an AI model, it required a complete rethinking of how data is collected, processed, and used across the network to support intelligent and automated decision-making. Building a dedicated platform to host data and develop intelligent applications was a critical first step toward solving this complex challenge. It provided the foundation for a sustainable and scalable approach to deploying AI in O-RAN environments. Such platforms are not readily available in the market and remain in the early stages of maturity, particularly for multi-vendor, production-grade telecom networks. Therefore, developing this platform was not just a support element, it was a necessary innovation that made it possible to manage complexity, enable automation, and ensure long-term adaptability of AI solutions in O-RAN. Conversely, relying on traditional data solutions from a single vendor is nearly impossible in a multi-vendor environment. These solutions are often limited in scope, requiring significant customization and additional tools to meet operational needs, ultimately leading to higher costs and complexity.
@@
Figure 1: Multi-vendor O-RAN network setup with Vendor-A in light blue, Vendor-B in green, and Vendor-Z in orange. In addition to other systems, equipment, and user devices that are working within the network and produce important data.
In this paper, we introduce a novel cloud-native, open data analytics platform for O-RAN to address these challenges and streamline the integration of AI applications into the O-RAN management stack. This platform leverages cutting-edge engineering technologies and concepts such as data lakehouse [7], DevOps [8], and MLOps. The platform is designed to support the growing number of private network deployments, regardless of the vendor, and the increasing complexity of their use cases. It is built to scale and handle the large amounts of data these networks generate. The data sources collected in the platform are big, variant and complex, coming from many different systems, as shown in Fig. 1. Leveraging this data, we applied a long short-term memory (LSTM)-based anomaly detection model to address the critical offshore connectivity challenge. By deploying the model in real time, the system enabled proactive anomaly detection and triggered automated mitigation actions, resulting in over 90% reduction in modem disconnections compared to the conventional method. This deployment represents one of the first real-world, production-grade demonstrations of AI-driven anomaly detection in a live, multi-vendor O-RAN network.
The main contributions of this work are summarized below and explained in more detail in the literature review (Sec. 2), where the novelty of our approach is also highlighted in relation to existing research.
@@
    To the best of the authors’ knowledge, we propose the first holistic, cloud-native data platform architecture for multi-vendor O-RAN environments, tightly aligned with O-RAN principles and designed to support scalable AI lifecycle integration. The platform also addresses gaps in standardization and engineering workflows for deploying AI in telecom networks.
@@
    The platform fully automates infrastructure setup, data handling, and AI pipelines using DevOps and GitOps technologies, significantly reducing operational workload and improving development efficiency.
@@
    We demonstrate the resolution of a real-time offshore connectivity problem in production using an LSTM-based anomaly detection model deployed within the platform, showcasing both technical feasibility and operational impact.
The rest of this paper is organized as follows. Sec. 2 introduces the background of O-RAN and how AI is being used to make networks more intelligent, along with a review of related work and the unique contributions and novelty of this study. Sec. 3 presents the real-world problem faced in offshore deployments and explains why AI was chosen to solve it and why the data platform was crucial to do that. Sect. 4 describes the design of the proposed cloud-native data platform, focusing on its architecture, key technologies, component layers, and how it supports standardizing the AI development. Sect. 5 details the AI-based anomaly detection use case, including model design, data preparation and feature engineering, deployment, and performance results. Sec. 6 discusses current limitations and future improvements, and Sec. 7 concludes the paper by summarizing the main findings and impact of this work.
2 Literature review: AI for O-RAN
In this section, we present the background information about the O-RAN technique, the relevant progress in research, and AI applications of O-RAN.
2.1 O-RAN background
The RAN is a critical component of a typical mobile communication network, enabling UE to connect to the core network, which then delivers services to users. The evolution of wireless communication systems from First-Generation (1G) to Fifth-Generation (5G) highlights increasing modularity and virtualization of network functionalities. Key advancements in RAN architecture include distributed RAN (D-RAN), centralized (or Cloud) RAN (C-RAN), and virtual RAN (vRAN). The distinctions among these architectures are detailed in [9].
@@ 2.1 O-RAN background
2.2 Literature review
Notably, recent advances in AI have led to increased exploration of its applications in communication networks, particularly through O-RAN architectures. Despite growing research interest, most studies are still based on simulations or lab setups. Deploying AI in real, production-grade O-RAN networks remains difficult because the technology is still new, and there are many technical and operational challenges to overcome.
For instance, Bonati et al. [13] introduced the OpenRAN Gym platform, which facilitates the development and testing of AI/ML algorithms on programmable RAN interfaces; however, their work is limited to testbed emulations rather than operational networks. Polese et al. [14] offered a comprehensive architectural and algorithmic survey of O-RAN’s intelligent functions, including the roles of RICs and xApps, but did not include any empirical system deployment. Similarly, Yeh et al. [15] and Nagib et al. [16] explored deep learning approaches for network slicing and traffic optimization within O-RAN settings, yet their implementations relied on idealized, offline datasets. Erdol et al. [17] proposed federated learning for traffic steering across distributed O-RANs, emphasizing privacy and coordination, but their framework has not been validated in real deployment scenarios. In [18], a deep reinforcement learning and graph neural network combined method is introduced to solve the baseband placement problem of O-RAN under different scenarios, but still it is based on a simulation environment.
Basaran and Dressler [19] addressed the explainability challenge through their XAInomaly framework, employing deep contractive autoencoders to detect anomalies in synthetic traffic data. While their focus on interpretability is valuable, the system lacks real-time capability and remains confined to simulated environments. Xavier et al. [20] tackled network security through supervised learning models for cyberattack detection in lab-based O-RAN testbeds. Despite achieving high accuracy, their study focuses solely on threat classification and not continuous anomaly detection in operational systems.
To date, no published studies have demonstrated the deployment of AI-driven models in live O-RAN environments using real network data, especially in business-critical settings like offshore networks, where connectivity issues can directly impact operations and safety. The present work addresses this critical gap by showcasing the design, implementation, and production deployment of an AI-enabled anomaly detection system within a live offshore mobile network built on O-RAN principles. Table 1 represents the most relevant recent research efforts and compares them with our contribution. What distinguishes this study is its integration of multiple advanced technologies such AI, big data processing, and cloud-native engineering within a multi-vendor, operational telecom infrastructure. A unified platform was developed in-house to support large-scale telemetry ingestion, real-time analytics, and continuous integration and deployment (CI/CD) workflows tailored specifically for the telemetry and control interfaces of O-RAN. Moreover, this solution uniquely incorporates live network data and UE feedback captured through vessel-based customer-premise equipment (CPE), offering a distinctive, UE perspective view of network performance and enabling faster, more accurate anomaly detection. This convergence of real-world deployment, cross-layer data fusion, and autonomous response positions the work as a novel and timely contribution to the field of intelligent and resilient mobile networks.
Table 1: Comparative Summary of Related Works
Study & Reference Deployment Environment Data Source Proposed AI-Data Platform AI Model Data Type Real-Time Key Outcomes
Ahmad et al., 2025 (This work) Production Offshore windfarms Network Real-world FCAPS & CPE telemetry Yes, (Cloud-native, multi-vendor) LSTM Network + CPE Yes 90% reduction in disconnections
Bonati et al., 2023 (OpenRAN Gym) Simulated testbed Synthetic data No Reinforcement Learning Network only Limited Demonstrated RIC optimization
Basaran & Dressler, 2025 (XAI anomaly) Simulated O-RAN Synthetic traffic data No Deep Contractive Autoencoder Network only No Introduced explainable anomaly detection
Xavier et al., 2023 (Early Detection) Lab testbed Simulated attack scenarios No ML (various) Network only Near real-time Achieved 95% accuracy in attack detection
2.3 Summary of the research novelty
This research presents a complete and practical innovation in the integration of AI and cloud-native architectures within real-world O-RAN deployments. The work begins with a rare and operationally critical business problem: maintaining stable connectivity in a private O-RAN network deployed across mobile offshore windfarm vessels. These networks face severe challenges, including unstable radio conditions, dynamic weather, signal interference over water, and vessel mobility. Unlike prior studies based on lab simulations or controlled testbeds, this research is rooted in an active offshore environment where system performance directly impacts business continuity.
To address this, a full-stack platform was designed and built using modern cloud-native engineering principles. The architecture integrates DevOps automation, Kubernetes orchestration, and scalable telemetry pipelines to support AI applications in production-grade settings. The platform supports data collection from both disaggregated O-RAN nodes and vessel-mounted customer equipment, enabling a unified view of network and user conditions. This dual-layer telemetry system is a rare feature in O-RAN AI literature and provides a foundation for context-aware diagnostics.
AI was added through a tightly integrated LSTM model pipeline trained to detect temporal anomalies in real time. The model’s ability to capture sequential patterns made it suitable for recognizing subtle faults in network behavior. Its integration with the platform enables anomaly detection and automated remediation within seconds, completing a closed-loop operational cycle that remains uncommon in O-RAN practice.
The platform was deployed and validated in a real offshore environment, achieving a reduction of over 90% in connectivity issues. This operational result demonstrates that AI can deliver a measurable impact in production disaggregated networks.
This work stands out not only for its use case and results, but also for delivering a clear architectural framework that supports standardization in AI deployment for O-RAN. It provides a blueprint for integrating automation, AI, and data engineering into complex, multi-vendor environments, where real-time performance and cross-layer visibility are essential.
By addressing both business and engineering challenges in a complete, deployable system, this research contributes a rare and practical advance in the evolving field of intelligent RANs.
3 Problem statement
This section starts by describing the real business problem that led to this work. We then explain why AI was seen as a good fit for solving problems in the context of O-RAN. After that, we discuss the main challenges we faced when trying to apply AI in such a complex, multi-vendor network. These points help show the necessity of building a dedicated data Analytics platform.
3.1 Business problem
Fig. 3 illustrates the business case of deploying a commercial O-RAN-based mobile private network (MPN) in an offshore location, serving the UEs that are located in vessels that navigate around the windfarm attending to wind turbines for infrastructure operational and maintenance activities. These vessels spend most of the time navigating close to the windfarm and carry tens of people who work at sea for periods of 15 days. Therefore, these people rely on this connectivity to do their work, to communicate with their colleagues working onshore, their families, and for their entertainment. The connectivity provided by this network also supports business and operational critical processes to the organizations responsible for operating and maintaining the network, across an area of around $300\,\text{km}^{2}$. Many factors such as weather conditions, sea conditions, distance to the site, and MPN operational faults may affect the quality of the connectivity service as experienced by the end-users, which might impact the: ability of running business and operational critical processes work according to the requirements or the ability of people working productively in their floating offices. Despite these challenges, the MPN’s network operator is responsible for managing, operating, and maintaining the network’s performance as per the contracted service level agreement.
Figure 3: Illustration for the real-time offshore 5G network deployment using O-RAN and challenges in managing coverage in the sea.
The problem addressed by the solution described in this paper occurred in an MPN service deployed by Boldyn Networks in a windfarm located in the North Sea. This network is composed of 3-cellular sites of 3 sectors each and one cellular carrier per sector, totaling 9 macro cells that cover the whole extension of the windfarm. The cellular sites are installed across the windfarm in three different turbines, where commercial-off-the-shelf (COTS) servers are used to run the containerized O-DU, and three O-RUs are installed to implement a sector each. The O-DUs deployed offshore connect, via dark-fibre light-up using long-haul SFPs, to an infrastructure of COTS servers that run the O-CU container. On the other hand, each relevant vessel’s network is composed of a Wi-Fi network that provides IP connectivity across the vessel. This Wi-Fi network connects to Boldyn’s customer CPE that acts like an LTE broadband router back-hauling the vessels’ IP traffic to connect to the internet and the MPN customer’s enterprise network. This CPE is installed inside the vessel, and it’s connected to external 4x4 multiple-input-multiple-output (MIMO) antennae that increase the coverage and capacity of the network. Each CPE is dual-modem capable for connection resilience and traffic load balancing reasons and one MPN SIM is inserted in each modem. The problem resides in the connection recovery procedure implemented by the CPE when one of the modems drops its link to the network. This procedure takes 5 minutes to re-establish a connection to the macro network, which is a long period of time and in some conditions increases the likelihood of disconnections occurring in both modems at the same time. In studying the problem, it was found the following phenomena:
  • It could be observed drift in the behavior of radio link performance, between the two modems of the CPE, verified by monitoring indicators such as latency and throughput, even if most of the time these were connected to the same cell and benefiting from similar radio conditions.
  • Updating the configuration of the modem on the CPE via its API, forced the modem connection to restart instantly and established a new radio link improving the performance of the modem in all the indicators.
  • The process of this type of restart in the modem is much quicker to implement, when the model is online, than the process of re-connection in case of radio link drop, and has the additional advantage that can be done pre-emptively whilst the other modem is showing degrading performance.
  • The automation of this process of analysis and decision-making could significantly reduce the number of disconnections, thus having a great impact on the quality of experience (QoE) of the users being served by the network system.
The automation of this process enabled us to develop a self-healing mechanism that implements advanced techniques. We tried conventional methods like rule-based systems, which rely on observed situations and the experience of the subject matter expert (SME). However, these methods often fall short in detecting disconnections before they occur and tend to perform many unnecessary actions, making them difficult to utilize and ineffective in handling complex and unexpected cases. In this case, using AI is considered the best approach to solve the problem, as it learns from patterns in historical data over time. For the detection of this type of performance anomaly in the modem, the model was trained on available historical performance data merged with records showing the timestamps when engineers took action to resolve the issue. This resulted in a labeled dataset that highlights the relationship between the behavior patterns of relevant performance indicators and the need for decision-making action. Some of these performance indicators included reference signal received power (RSRP), reference signal received quality (RSRQ), IP packet data latency, timestamp, location, the cell to which the modem is connected, and others from various parts of the network.
The process of collecting data from the CPE device via the API every 5 seconds, preparing and enriching it, and streaming it to the proposed data platform was designed to enable real-time decision-making and direct provisioning through machine learning. In this offshore environment, where disruptions must be detected and resolved within seconds, AI plays a central role in maintaining service continuity and operational efficiency. However, this intelligence can only operate effectively when supported by a robust data platform. The platform provides the necessary infrastructure to manage, transform, and deliver high-frequency telemetry data to the model in a reliable and scalable manner. Without it, the AI pipeline would lack the consistency, speed, and data quality required for low-latency inference and autonomous action. Thus, the data management platform is not just a support layer but a critical enabler of the AI-driven solution.
@@ 3.2 The motivation of deploying AI in O-RAN
  • Reducing Complexity: O-RAN networks have a more complex, disaggregated architecture compared to traditional RAN, making manual management and optimization more challenging. Also, building traditional applications that process the data will be challenging. AI algorithms can automate and optimize these processes, like the discussed auto-healing process in this paper. It also compensates for the shortage of skilled engineers to manage such novel networks  [21].
@@
    • Improved Network Performance: AI-based algorithms can be used to optimize network performance by dynamically allocating network resources and adjusting network parameters based on real-time network conditions such as autonomous QoE and QoS resource optimization [4, 21, 23].